Shift report 20020115 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Jan 15 2002 - 23:54:29 EST

  • Next message: Apache: "Shift report 20020116 00:00-08:00"

    ------------------------------------------------------------
    Date: Tue Jan 15 23:54:29 2002, Shift: 16:00-24:00
    Supervisor: Mads, with: JH, Ramiro
    Log book#12 pages 65 - 66
    
    Shift Summary:
    ------------------------------------------------------------
    Ramiro and JH seem to have soleved the problem with T2 that has disabled us to take data from the FFS the previous 2 shifts. We have been taking some data during the last 4 hours of this shift: 6593,6594,6596  
    
    The magnet field for D1-D4 was changed to 0A, and later to 1/3B. 
    
    For some reason we cold not move the back FFS when we atempted to do so. 
    
    There has also been problems whit T3-T5 tripping several times. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Jan 15 2002 - 23:54:57 EST